% Generated by IEEEtran.bst, version: 1.14 (2015/08/26)
\begin{thebibliography}{10}
\providecommand{\url}[1]{#1}
\csname url@samestyle\endcsname
\providecommand{\newblock}{\relax}
\providecommand{\bibinfo}[2]{#2}
\providecommand{\BIBentrySTDinterwordspacing}{\spaceskip=0pt\relax}
\providecommand{\BIBentryALTinterwordstretchfactor}{4}
\providecommand{\BIBentryALTinterwordspacing}{\spaceskip=\fontdimen2\font plus
\BIBentryALTinterwordstretchfactor\fontdimen3\font minus
  \fontdimen4\font\relax}
\providecommand{\BIBforeignlanguage}[2]{{%
\expandafter\ifx\csname l@#1\endcsname\relax
\typeout{** WARNING: IEEEtran.bst: No hyphenation pattern has been}%
\typeout{** loaded for the language `#1'. Using the pattern for}%
\typeout{** the default language instead.}%
\else
\language=\csname l@#1\endcsname
\fi
#2}}
\providecommand{\BIBdecl}{\relax}
\BIBdecl

\bibitem{SharmaS20}
\BIBentryALTinterwordspacing
G.~G. Sharma and K.~Stol, ``Exploring onboarding success, organizational fit,
  and turnover intention of software professionals,'' \emph{J. Syst. Softw.},
  vol. 159, 2020. [Online]. Available:
  \url{https://doi.org/10.1016/j.jss.2019.110442}
\BIBentrySTDinterwordspacing

\bibitem{Steinmacher15}
\BIBentryALTinterwordspacing
I.~Steinmacher, M.~A.~G. Silva, M.~A. Gerosa, and D.~F. Redmiles, ``A
  systematic literature review on the barriers faced by newcomers to open
  source software projects,'' \emph{Inf. Softw. Technol.}, vol.~59, pp. 67--85,
  2015. [Online]. Available: \url{https://doi.org/10.1016/j.infsof.2014.11.001}
\BIBentrySTDinterwordspacing

\bibitem{Balali18}
\BIBentryALTinterwordspacing
S.~Balali, I.~Steinmacher, U.~Annamalai, A.~Sarma, and M.~A. Gerosa,
  ``Newcomers' barriers. . . is that all? an analysis of mentors' and
  newcomers' barriers in {OSS} projects,'' \emph{Comput. Support. Cooperative
  Work.}, vol.~27, no. 3-6, pp. 679--714, 2018. [Online]. Available:
  \url{https://doi.org/10.1007/s10606-018-9310-8}
\BIBentrySTDinterwordspacing

\bibitem{RastogiT0NC17}
\BIBentryALTinterwordspacing
A.~Rastogi, S.~Thummalapenta, T.~Zimmermann, N.~Nagappan, and J.~Czerwonka,
  ``Ramp-up journey of new hires: Do strategic practices of software companies
  influence productivity?'' in \emph{Proceedings of the 10th Innovations in
  Software Engineering Conference, {ISEC} 2017, Jaipur, India, February 5-7,
  2017}.\hskip 1em plus 0.5em minus 0.4em\relax {ACM}, 2017, pp. 107--111.
  [Online]. Available: \url{http://dl.acm.org/citation.cfm?id=3021471}
\BIBentrySTDinterwordspacing

\bibitem{bauer11}
T.~N. Bauer and B.~Erdogan, ``{Organizational socialization: The effective
  onboarding of new employees.}'' Washington, DC, US, pp. 51--64, 2011.

\bibitem{Viviani19}
\BIBentryALTinterwordspacing
G.~Viviani and G.~C. Murphy, ``Reflections on onboarding practices in mid-sized
  companies,'' in \emph{Proceedings of the 12th International Workshop on
  Cooperative and Human Aspects of Software Engineering, CHASE@ICSE 2019,
  Montr{\'{e}}al, QC, Canada, 27 May 2019}.\hskip 1em plus 0.5em minus
  0.4em\relax {IEEE} / {ACM}, 2019, pp. 83--84. [Online]. Available:
  \url{https://doi.org/10.1109/CHASE.2019.00027}
\BIBentrySTDinterwordspacing

\bibitem{Clements02}
P.~Clements and L.~M. Northrop, \emph{Software product lines - practices and
  patterns}, ser. {SEI} series in software engineering.\hskip 1em plus 0.5em
  minus 0.4em\relax Addison-Wesley, 2002.

\bibitem{Deelstra05}
\BIBentryALTinterwordspacing
S.~Deelstra, M.~Sinnema, and J.~Bosch, ``Product derivation in software product
  families: a case study,'' \emph{J. Syst. Softw.}, vol.~74, no.~2, pp.
  173--194, 2005. [Online]. Available:
  \url{https://doi.org/10.1016/j.jss.2003.11.012}
\BIBentrySTDinterwordspacing

\bibitem{Kolb06}
\BIBentryALTinterwordspacing
R.~Kolb, I.~John, J.~Knodel, D.~Muthig, U.~Haury, and G.~Meier, ``Experiences
  with product line development of embedded systems at testo {AG},'' in
  \emph{Software Product Lines, 10th International Conference, {SPLC} 2006,
  Baltimore, Maryland, USA, August 21-24, 2006, Proceedings}.\hskip 1em plus
  0.5em minus 0.4em\relax {IEEE} Computer Society, 2006, pp. 172--181.
  [Online]. Available: \url{https://doi.org/10.1109/SPLINE.2006.1691589}
\BIBentrySTDinterwordspacing

\bibitem{Deelstra04}
\BIBentryALTinterwordspacing
S.~Deelstra, M.~Sinnema, and J.~Bosch, ``Experiences in software product
  families: Problems and issues during product derivation,'' in \emph{Software
  Product Lines, Third International Conference, {SPLC} 2004, Boston, MA, USA,
  August 30-September 2, 2004, Proceedings}, ser. Lecture Notes in Computer
  Science, R.~L. Nord, Ed., vol. 3154.\hskip 1em plus 0.5em minus 0.4em\relax
  Springer, 2004, pp. 165--182. [Online]. Available:
  \url{https://doi.org/10.1007/978-3-540-28630-1\_10}
\BIBentrySTDinterwordspacing

\bibitem{Capilla13}
R.~Capilla, J.~Bosch, and K.-C. Kang, \emph{Systems and Software Variability
  Management: Concepts, Tools and Experiences}.\hskip 1em plus 0.5em minus
  0.4em\relax Springer Publishing Company, Incorporated, 2013.

\bibitem{Ernst02}
\BIBentryALTinterwordspacing
M.~D. Ernst, G.~J. Badros, and D.~Notkin, ``An empirical analysis of {C}
  preprocessor use,'' \emph{{IEEE} Trans. Software Eng.}, vol.~28, no.~12, pp.
  1146--1170, 2002. [Online]. Available:
  \url{https://doi.org/10.1109/TSE.2002.1158288}
\BIBentrySTDinterwordspacing

\bibitem{Melo17}
\BIBentryALTinterwordspacing
J.~Melo, F.~B. Narcizo, D.~W. Hansen, C.~Brabrand, and A.~Wasowski,
  ``Variability through the eyes of the programmer,'' in \emph{Proceedings of
  the 25th International Conference on Program Comprehension, {ICPC} 2017,
  Buenos Aires, Argentina, May 22-23, 2017}.\hskip 1em plus 0.5em minus
  0.4em\relax {IEEE} Computer Society, 2017, pp. 34--44. [Online]. Available:
  \url{https://doi.org/10.1109/ICPC.2017.34}
\BIBentrySTDinterwordspacing

\bibitem{Acher17}
\BIBentryALTinterwordspacing
M.~Acher, R.~E. Lopez{-}Herrejon, and R.~Rabiser, ``Teaching software product
  lines: {A} snapshot of current practices and challenges,'' \emph{{ACM}
  Transactions on Computer Education}, vol.~18, no.~1, pp. 2:1--2:31, 2017.
  [Online]. Available: \url{https://doi.org/10.1145/3088440}
\BIBentrySTDinterwordspacing

\bibitem{quintana2004scaffolding}
C.~Quintana, B.~J. Reiser, E.~A. Davis, J.~Krajcik, E.~Fretz, R.~G. Duncan,
  E.~Kyza, D.~Edelson, and E.~Soloway, ``A scaffolding design framework for
  software to support science inquiry,'' \emph{The journal of the learning
  sciences}, vol.~13, no.~3, pp. 337--386, 2004.

\bibitem{Heimburger20}
L.~Heimburger, L.~Buchweitz, R.~Gouveia, and O.~Korn, ``Gamifying onboarding:
  How to increase both engagement and integration of new employees,'' in
  \emph{Advances in Social and Occupational Ergonomics}.\hskip 1em plus 0.5em
  minus 0.4em\relax Cham: Springer International Publishing, 2020, pp. 3--14.

\bibitem{Hibschman19}
\BIBentryALTinterwordspacing
J.~Hibschman, D.~Gergle, E.~O'Rourke, and H.~Zhang, ``Isopleth: Supporting
  sensemaking of professional web applications to create readily available
  learning experiences,'' \emph{{ACM} Trans. Comput. Hum. Interact.}, vol.~26,
  no.~3, pp. 16:1--16:42, 2019. [Online]. Available:
  \url{https://doi.org/10.1145/3310274}
\BIBentrySTDinterwordspacing

\bibitem{CubranicM03}
\BIBentryALTinterwordspacing
D.~Cubranic and G.~C. Murphy, ``Hipikat: Recommending pertinent software
  development artifacts,'' in \emph{Proceedings of the 25th International
  Conference on Software Engineering, May 3-10, 2003, Portland, Oregon,
  {USA}}.\hskip 1em plus 0.5em minus 0.4em\relax {IEEE} Computer Society, 2003,
  pp. 408--418. [Online]. Available:
  \url{https://doi.org/10.1109/ICSE.2003.1201219}
\BIBentrySTDinterwordspacing

\bibitem{Panichella15}
\BIBentryALTinterwordspacing
S.~Panichella, ``Supporting newcomers in software development projects,'' in
  \emph{2015 {IEEE} International Conference on Software Maintenance and
  Evolution, {ICSME} 2015, Bremen, Germany, September 29 - October 1,
  2015}.\hskip 1em plus 0.5em minus 0.4em\relax {IEEE} Computer Society, 2015,
  pp. 586--589. [Online]. Available:
  \url{https://doi.org/10.1109/ICSM.2015.7332519}
\BIBentrySTDinterwordspacing

\bibitem{MalheirosMTM12}
\BIBentryALTinterwordspacing
Y.~Malheiros, A.~Moraes, C.~Trindade, and S.~Meira, ``A source code recommender
  system to support newcomers,'' in \emph{36th Annual {IEEE} Computer Software
  and Applications Conference, {COMPSAC} 2012, Izmir, Turkey, July 16-20,
  2012}.\hskip 1em plus 0.5em minus 0.4em\relax {IEEE} Computer Society, 2012,
  pp. 19--24. [Online]. Available:
  \url{https://doi.org/10.1109/COMPSAC.2012.11}
\BIBentrySTDinterwordspacing

\bibitem{WangS11}
\BIBentryALTinterwordspacing
J.~Wang and A.~Sarma, ``Which bug should {I} fix: helping new developers
  onboard a new project,'' in \emph{Proceedings of the 4th International
  Workshop on Cooperative and Human Aspects of Software Engineering, {CHASE}
  2011, Waikiki, Honolulu, HI, USA, May 21, 2011}.\hskip 1em plus 0.5em minus
  0.4em\relax {ACM}, 2011, pp. 76--79. [Online]. Available:
  \url{https://doi.org/10.1145/1984642.1984661}
\BIBentrySTDinterwordspacing

\bibitem{SteinmacherCTG16}
\BIBentryALTinterwordspacing
I.~Steinmacher, T.~U. Conte, C.~Treude, and M.~A. Gerosa, ``Overcoming open
  source project entry barriers with a portal for newcomers,'' in
  \emph{Proceedings of the 38th International Conference on Software
  Engineering, {ICSE} 2016, Austin, TX, USA, May 14-22, 2016}.\hskip 1em plus
  0.5em minus 0.4em\relax {ACM}, 2016, pp. 273--284. [Online]. Available:
  \url{https://doi.org/10.1145/2884781.2884806}
\BIBentrySTDinterwordspacing

\bibitem{JohannessonP14}
\BIBentryALTinterwordspacing
P.~Johannesson and E.~Perjons, \emph{An Introduction to Design Science}.\hskip
  1em plus 0.5em minus 0.4em\relax Springer, 2014. [Online]. Available:
  \url{https://doi.org/10.1007/978-3-319-10632-8}
\BIBentrySTDinterwordspacing

\bibitem{Apel13}
\BIBentryALTinterwordspacing
S.~Apel, D.~S. Batory, C.~K{\"{a}}stner, and G.~Saake, \emph{Feature-Oriented
  Software Product Lines - Concepts and Implementation}.\hskip 1em plus 0.5em
  minus 0.4em\relax Springer, 2013. [Online]. Available:
  \url{https://doi.org/10.1007/978-3-642-37521-7}
\BIBentrySTDinterwordspacing

\bibitem{AcherLR18}
\BIBentryALTinterwordspacing
M.~Acher, R.~E. Lopez{-}Herrejon, and R.~Rabiser, ``Teaching software product
  lines: a snapshot of current practices and challenges (journal-first
  abstract),'' in \emph{Proceeedings of the 22nd International Systems and
  Software Product Line Conference - Volume 1, {SPLC} 2018, Gothenburg, Sweden,
  September 10-14, 2018}.\hskip 1em plus 0.5em minus 0.4em\relax {ACM}, 2018,
  p. 249. [Online]. Available: \url{https://doi.org/10.1145/3233027.3236393}
\BIBentrySTDinterwordspacing

\bibitem{novak2006theory}
J.~D. Novak and A.~J. Ca{\~n}as, ``The theory underlying concept maps and how
  to construct them,'' \emph{Florida Institute for Human and Machine
  Cognition}, vol.~1, no.~1, pp. 1--31, 2006.

\bibitem{nesbit2006learning}
J.~C. Nesbit and O.~O. Adesope, ``Learning with concept and knowledge maps: A
  meta-analysis,'' \emph{Review of educational research}, vol.~76, no.~3, pp.
  413--448, 2006.

\bibitem{AbilhoaC14}
\BIBentryALTinterwordspacing
W.~D. Abilhoa and L.~N. de~Castro, ``A keyword extraction method from twitter
  messages represented as graphs,'' \emph{Appl. Math. Comput.}, vol. 240, pp.
  308--325, 2014. [Online]. Available:
  \url{https://doi.org/10.1016/j.amc.2014.04.090}
\BIBentrySTDinterwordspacing

\bibitem{Wohlin12}
\BIBentryALTinterwordspacing
C.~Wohlin, P.~Runeson, M.~H{\"{o}}st, M.~C. Ohlsson, and B.~Regnell,
  \emph{Experimentation in Software Engineering}.\hskip 1em plus 0.5em minus
  0.4em\relax Springer, 2012. [Online]. Available:
  \url{https://doi.org/10.1007/978-3-642-29044-2}
\BIBentrySTDinterwordspacing

\bibitem{Berger15}
\BIBentryALTinterwordspacing
T.~Berger, D.~Lettner, J.~Rubin, P.~Gr{\"{u}}nbacher, A.~Silva, M.~Becker,
  M.~Chechik, and K.~Czarnecki, ``What is a feature?: a qualitative study of
  features in industrial software product lines,'' in \emph{Proceedings of the
  19th International Conference on Software Product Line, {SPLC} 2015,
  Nashville, TN, USA, July 20-24, 2015}, D.~C. Schmidt, Ed.\hskip 1em plus
  0.5em minus 0.4em\relax {ACM}, 2015, pp. 16--25. [Online]. Available:
  \url{https://doi.org/10.1145/2791060.2791108}
\BIBentrySTDinterwordspacing

\bibitem{Begel08}
\BIBentryALTinterwordspacing
A.~Begel and B.~Simon, ``Struggles of new college graduates in their first
  software development job,'' in \emph{Proceedings of the 39th {SIGCSE}
  Technical Symposium on Computer Science Education, {SIGCSE} 2008, Portland,
  OR, USA, March 12-15, 2008}.\hskip 1em plus 0.5em minus 0.4em\relax {ACM},
  2008, pp. 226--230. [Online]. Available:
  \url{https://doi.org/10.1145/1352135.1352218}
\BIBentrySTDinterwordspacing

\bibitem{Detienne01}
\BIBentryALTinterwordspacing
F.~D{\'{e}}tienne, \emph{Software design cognitive aspects}, ser. Practitioner
  series.\hskip 1em plus 0.5em minus 0.4em\relax Springer, 2001. [Online].
  Available: \url{http://www.springer.com/computer/swe/book/978-1-85233-253-2}
\BIBentrySTDinterwordspacing

\bibitem{Runeson12}
\BIBentryALTinterwordspacing
P.~Runeson, M.~H{\"{o}}st, A.~Rainer, and B.~Regnell, \emph{Case Study Research
  in Software Engineering - Guidelines and Examples}.\hskip 1em plus 0.5em
  minus 0.4em\relax Wiley, 2012. [Online]. Available:
  \url{http://eu.wiley.com/WileyCDA/WileyTitle/productCd-1118104358.html}
\BIBentrySTDinterwordspacing

\end{thebibliography}

\documentclass[10pt,conference]{IEEEtran}
\IEEEoverridecommandlockouts
\usepackage{cite}
\usepackage{amsmath,amssymb,amsfonts}
\usepackage{algorithmic}
\usepackage{graphicx}
\usepackage{comment}
\usepackage{textcomp}
\usepackage{xcolor}
\def\BibTeX{{\rm B\kern-.05em{\sc i\kern-.025em b}\kern-.08em
    T\kern-.1667em\lower.7ex\hbox{E}\kern-.125emX}}
\begin{document}
\title{Unburdening onboarding in Software Product Lines}
\author{\IEEEauthorblockN{Raul Medeiros}
\IEEEauthorblockA{University of the Basque Country (UPV/EHU)\\
San Sebasti{\'a}n, Spain\\
raul.medeiros@ehu.eus}
}
\maketitle
\renewcommand\IEEEkeywordsname{Keywords}

\begin{abstract}
The number of studies focusing on onboarding in software organizations has increased significantly during the last years. However, current literature overlooks onboarding in Software Product Lines (SPLs). SPLs have been proven effective in managing the increasing variability of industry software and enabling systematic reuse of a product family. Despite these benefits, SPLs are complex and exhibit particular characteristics that distinguish them from traditional software. Due to these peculiarities, SPLs require a tailor-made onboarding process. Assistance tools might help. In this dissertation, we propose assistance tools (i.e., tools built on top of the software project that help  learners  understand and develop  knowledge) as a means for helping newcomers during onboarding in SPLs.
\end{abstract}
\begin{IEEEkeywords}
Software Product Lines,
Onboarding,
Recommender Systems,
Concept Maps
\end{IEEEkeywords}
\section{Introduction}\label{intro}
Today's world is highly interconnected. In such a situation, professionals often move from one company to another, especially inside the software industry\cite{SharmaS20}. High mobility implies new hired developers, who must be introduced into the company's culture, processes, technologies, etc., by means of the fastest and most effective possible approach. In the literature, this introduction process is known as \textit{onboarding}\cite{Steinmacher15,Balali18,RastogiT0NC17,bauer11}. Yet, incorporating into a new software development team is not a trivial task, and it has been studied that newcomers (i.e., people incorporating to the company) face several barriers during onboarding\cite{Balali18}, namely:

\begin{itemize}
    \item \textit{technical barriers}, newcomers often encounter themselves without knowing how to set up their development environment or lack the skills to manage the programming languages used in the organization. This problem is accompanied by the lack of prior knowledge of the domain where development takes place. 
    \item \textit{process barriers},  software projects usually lack a formal procedure for introducing newcomers.
    \item \textit{interpersonal barriers}, which refer to the absence of communication and socializing skills in the newcomer. These barriers arise more fragrantly when newcomers are incorporated into a diverse team, where different people with different interpersonal skills gather. 
    \item \textit{personal barriers} include issues related to personality, lack of management skills, or behavior problems. 
\end{itemize}

One of the used approaches to alleviate these barriers consists in considering onboarding as a mentor-newcomer journey (aka mentoring), where the senior developer (i.e., the mentor) assists the newcomer (i.e., the mentee) in her/his incorporation,  both transferring technical skills and knowledge, and providing moral support\cite{SharmaS20}. However, senior developers' time is a valued resource and spending it teaching newcomers reduces significantly their productivity. Therefore, most of the companies expect newcomers to explore and understand the software by themselves\cite{Viviani19}.

Onboarding barriers exacerbate in Software Product Lines (SPL), which are both complex organizationally and software wise. The main goal of SPLs is to support the reuse of a whole family of software products in a systematic way\cite{Clements02}. Specifically, SPLs aim at automating software derivation (i.e., products of the SPL) out of a set of reusable assets (i.e., the SPL platform). Ideally, software derivation is limited to indicating the set of features to be exhibited by the desired products (so-called ``product configuration''). Features are user-visible characteristics of a software system.  In this scenario, products are obtained through a fully ``Configurable Product Family'' where product customization does not exist \cite{Deelstra05}. However, reaching a ``Configurable Product Family'' is estimated to last around ten years\cite{Kolb06}. In this transition, not only traditional onboarding occurs, but also developers are transferred from Application Engineering (AE) to Domain Engineering (DE) (aka \textbf{crossboarding})\cite{Deelstra04}. From now on, we refer to these developers transitioning from AE to DE as \textbf{crosscomers}.

SPLs pose unique challenges to crosscomers, setting the large volume of a SPL aside, crosscomers must adapt themselves to variability and  developing software for reuse. First, feature models gather together multiple concepts that are unknown for the crosscomer, leading to an increase in complexity and making them difficult to comprehend or even maintain\cite{Capilla13}. Moreover, the code should account for reuse. Often, this is achieved by enclosing variants within \#ifdef and \#endif directives, and associated with features. As a result, code comprehensibility is diminished \cite{Ernst02,Melo17}. In summary, the main barriers a crosscomer face when incorporating to the domain engineering team are the sheer  volume,  understanding  variability  and  internalizing  development  for  reuse  \cite{Acher17}.


This thesis focuses on the above-mentioned  barriers. To this end, we resort to assistance tools, which are tools that can help learners understand and develop knowledge (e.g., sensemaking scaffolds\cite{quintana2004scaffolding}). In this context, we pose our research hypothesis as follows: \textbf{\textit{Assistance tools help crosscomers during crossboarding in SPL organizations, simplifying and enhancing the process of incorporating to a domain engineering team}}.

To evaluate the hypothesis, we ask the following research
questions:

\begin{itemize}
    \item \textbf{RQ1:} How can sense-making scaffolds construct domain knowledge on top of crosscomers' background? \textbf{Study-1}
    \item \textbf{RQ2:} What is an appropriate journey for crosscomers to understand variability? \textbf{Study-2}
\end{itemize}
\section{Related Work}
To the best of our knowledge, there is no prior work that addresses onboarding or crossboarding  in SPLs. However, easing and speeding up onboarding has been the subject of distinct approaches:  setting a gamification system \cite{Heimburger20}; tools for assisted coding  \cite{Hibschman19}; visualization of data and conceptual representations \cite{Hibschman19}; recommendation assistants \cite{CubranicM03, Panichella15, MalheirosMTM12,WangS11}; setting a web portal\cite{SteinmacherCTG16}. Specifically, Malheiros et al. \cite{MalheirosMTM12} and Cubranic et al. \cite{CubranicM03} present two tools that support newcomers by recommending the most appropriate source code files for their development tasks. Wang and Sarma \cite{WangS11} focus on bug fixing recommendations, and they provide newcomers a term-based search feature with which newcomers can find bugs on a specific topic. Dominic et al. The \textit{Isopleth} tool \cite{Hibschman19}, the closest approach to this thesis, is a web-based platform that helps newcomers make sense of complex web software projects.


\section{Expected contributions and evaluation plans}

This thesis follows a Design Science Research (DSR) approach. DSR is the scientific study and creation of artifacts as they are developed and used by people with the goal of solving practical problems of general interest\cite{JohannessonP14}. Therefore, the contributions of this thesis have an artifact associated with it. Next, we delve into the proposed studies:

\subsection{Study-1: Domain knowledge construction}
\textbf{\textit{Problem.}} When crosscomers are incorporated into the domain engineering team, they have to deal with both variability and development for reuse. At this point, it is of critical importance that crosscomers understand the concepts and terminology used by the practicioners and how the SPL products are built \cite{Apel13}. However, directly exploring the SPL documentation (e.g. the feature model) or code to acquire domain knowledge is not feasible for crosscomers due to the sheer volume and variability of the SPL\cite{AcherLR18}\cite{Capilla13}. 
\textbf{\textit{Approach.}} In this study, we focus on helping crosscomers understand the set of concepts and terminology of the domain of the SPL platform. To this end, we propose the use of sensemaking scaffolds for presenting such data\cite{quintana2004scaffolding} and specifically concept maps as the realization of such scaffolds. A concept map is a diagram that depicts suggested relationships between concepts\cite{novak2006theory}. 
Concept mapping is reckoned to be a means for meaningful learning insofar as it serves as a kind of scaffold to help to organize knowledge\cite{nesbit2006learning}. Particularly, SPL concept maps gather domain concepts extracted from all available artifacts of the SPL. In our approach, features drive the extraction process. For each feature, its related concepts are extracted from SPL artifacts using NLP techniques such as keyword extraction\cite{AbilhoaC14}. Then, each concept receives a relevance. Finally, an expert selects the most representative concepts of the SPL and creates relationships between them completing the concept map.
\textbf{\textit{Evaluation plan.}} We want to evaluate the effect of concept maps in the acquisition of domain knowledge. This will be carried out with a controlled experiment, with one of our industrial partners, following the experiment design and reporting guidelines proposed by Wohlin et al. \cite{Wohlin12}. To this end, the evaluation will compare the domain knowledge of crosscomers using concept maps with the domain knowledge of those using other kinds of domain representations (e.g. feature models or entity-relationship diagrams), after having performed the given tasks.

\subsection{Study-2: Crossboarding journey}
\textbf{\textit{Problem.}} During crossboarding, crosscomers must move from development with reuse to development for reuse. This means that they have to deal with variability. In SPLs, variability is characterized in terms of  functional abstractions that serve for communicating, reasoning and distinguishing among individual products (i.e. features)\cite{Berger15}. Preprocessors are often the means to implement the features, and has been studied that they hinder program comprehension\cite{Ernst02,Melo17}. Understanding features can be a hard work and it is being reported that onboarding plans that are not structured to newcomers' knowledge, lead to newcomers' frustration\cite{Begel08}. This can also occur in crossboarding and, therefore, crosscomers need to be assigned with features that resemble to their prior knowledge as application engineers. 
\textbf{\textit{Approach.}} In crossboarding, the crosscomer profile can be characterized in terms of the codebase they have been in charge of (i.e. the code they evolved during their time in AE). We refer to this codebase as ``the Background Feature''. Akin to program comprehension theory, we propose to construct crosscomers' journeys (i.e. the order in which they should understand features) based on the similarity between the features and crosscomers' ``Background Feature''\cite{Detienne01}. To define these journeys, we resort to  Recommender Systems (RSs) based on feature similarity. Using Topic Modeling, the RS will create a similarity matrix, which can be represented as a weighted graph. Using this similarity graph, the recommender system will propose the easiest onboarding journey (i.e., a subgraph) based on similarity degrees.
\textbf{\textit{Evaluation plan.}} We plan to evaluate the recommended crossboarding journeys through a case study in one of our industrial partners or in WacLine our academic SPL\footnote{https://onekin.github.io/WacLine/}. The case study will follow Runeson et
al. \cite{Runeson12} guidelines and will try to identify the acceptation and
benefits of the crossboarding journeys in a real SPL crossboarding setting with at least four crosscomers. 


\section*{Acknowledgments}
This work is supported by Spanish Ministry of Science and Innovation (RTI2018-099818-B-I00) and the Ministry of Education (MCIU-AEI TIN2017-90644-REDT (TASOVA)). R. Medeiros enjoys a doctoral grant from the Ministry of
Science and Innovation.
\bibliographystyle{IEEEtran}
\bibliography{20SymposiumICSE}

\end{document}